So I have been reading this question for the past 37 minutes, and I cannot think of the equation.
All I need help on is setting up the starting equation....I know how to solve it from there.
If you could show me/help me set up the starting equation, I would be VERY grateful.

Terry has walked 3 miles. He averages 4 mi/hr. In how many more hours will he have walked 13 miles?

Found 2 solutions by Gmoney, Earlsdon:
Answer by Gmoney(2)

Well D= R T is a famous equation used to solve or write equations. It stands for distance=rate times time.
D=r t


Answer by Earlsdon(6294)

Since Terry has already walked 3 of the 13 miles, you know that he has (13-3) miles to go.
Remember that: d = rt or distance = rate(speed) multiplied by time.
You can write this as:
t = d/r and in this problem, r = 4 mph and d = (13-3), so you can write the equation:
t = (13-3)miles/4mph This will tell you how long, in hours, it will take Terry to cover the remaining miles.